# Insert-Only versus Insert-Delete in Dynamic Query Evaluation

### v3

@@ Abstract.
We study the dynamic query evaluation problem: Given a join query $Q$ and a sequence of updates, we would like to construct a data structure that supports constant-delay enumeration of the query output after each update.
We show that a sequence of $N$ insert-only updates to an initially empty database can be executed in total time $\mathcal{O}(N^{\mathsf{w}(Q)})$, where $\mathsf{w}(Q)$ is the fractional hypertree width of $Q$. This matches the complexity of the static query evaluation problem for $Q$ and a database of size $N$. One corollary is that the amortized time per single-tuple insert is constant for acyclic joins.
In contrast, we show that a sequence of $N$ inserts and deletes to $Q$ can be executed in total time $\widetilde{\mathcal{O}}(N^{\mathsf{w}(\widehat{Q})})$, where $\widehat{Q}$ is obtained from $Q$ by extending every relational atom with extra variables that represent the “lifespans” of tuples in $Q$. We show that this reduction is optimal in the sense that the static evaluation runtime of $\widehat{Q}$ provides a lower bound on the total update time of $Q$. Our approach recovers the optimal single-tuple update time for known queries such as the hierarchical and Loomis-Whitney join queries.
@@ 1. Introduction
Answering queries under updates to the underlying database, also called dynamic query evaluation, is a fundamental problem in data management, with recent work addressing it from both systems and theoretical perspectives. There have been recent efforts on building systems for dynamic query evaluation, such as DBToaster (Koch et al., 2014), DynYannakakis (Idris et al., 2017, 2020), F-IVM (Nikolic and Olteanu, 2018; Nikolic et al., 2020; Kara et al., 2023b), and CROWN (Wang et al., 2023). By allowing tuples to carry payloads with elements from rings (Koch, 2010), such systems can maintain complex analytics over database queries, such as linear algebra computation (Nikolic et al., 2014), collection programming (Koch et al., 2016), and machine learning models (Kara et al., 2023b). Beyond conjunctive queries, there has been work on intersection joins under updates (Tao and Yi, 2022), and dynamic computation for expressive languages such as Datalog (Motik et al., 2019), Differential Datalog (Abadi et al., 2015), or DBSP (Budiu et al., 2023). The descriptive complexity of various recursive Datalog queries under updates, such as reachability, has also been investigated (Datta et al., 2015, 2018; Schwentick et al., 2020).
@@
While this problem attracted continuous interest in academia and industry over the past decades, it is only relatively recently that the first results on the fine-grained complexity and in particular on the optimality of this problem have emerged. These efforts aim to mirror the breakthrough made by the introduction of worst-case optimal join algorithms in the database theory literature (Atserias et al., 2013; Veldhuizen, 2014; Ngo et al., 2018). Beyond notable yet limited explorations, understanding the optimality of maintenance for the entire language of conjunctive queries remains open. Prime examples of progress towards the optimality of query maintenance are the characterizations of queries that admit (worst-case or amortized) constant time per single-tuple update (insert or delete): the $q$-hierarchical queries (Berkholz et al., 2017; Idris et al., 2017; Kara et al., 2023b; Wang et al., 2023), queries that become $q$-hierarchical in the presence of free access patterns (Kara et al., 2023a) or by rewriting under functional dependencies (Kara et al., 2023b), or queries on update sequences whose enclosureness is bounded by a constant (Wang and Yi, 2020). The $\delta_{1}$-hierarchical queries (Kara et al., 2020b, 2023c, 2023a) and the triangle queries (Kara et al., 2019, 2020a) admit optimal, albeit non-constant, update time conditioned on the Online Matrix-Vector Multiplication (OMv) conjecture (Henzinger et al., 2015).
In this paper, we complement the above work in five ways.
@@
First, we give reductions from the dynamic query evaluation problem to the static query evaluation problem and vice versa. These reductions use a translation of the time dimension in the dynamic problem into a spatial dimension in the static problem, so that the maintenance under a stream of updates corresponds to taking the intersection of intervals representing the “lifespans” of tuples in the update stream. Our reductions link the complexities of the dynamic and static problems and allow to transfer both algorithms and lower bounds from one problem to the other.
Second, we pinpoint the complexity gap between the insert-only and the insert-delete settings in the dynamic evaluation of any (full conjunctive) query $Q$. Given a stream of updates to an initially empty database, our goal is to execute these updates one by one, while supporting constant-delay enumeration of the query result after each update. In the insert-only setting, we show that a stream of inserts to $Q$ can be executed one insert at a time in the same overall time as if the entire stream was executed as one bulk update (which corresponds to static evaluation). In contrast, in the insert-delete setting, a stream of both inserts and deletes to $Q$ can be executed one update at a time in the same overall time as the static evaluation of another query $\widehat{Q}$, which is obtained by extending every relational atom in $Q$ with extra variables representing the “lifespans” of tuples of $Q$. This query $\widehat{Q}$ may have a higher complexity than $Q$. Our reduction from the maintenance of $Q$ to the static evaluation problem for $\widehat{Q}$ is nevertheless optimal: Lower bounds on the latter problem imply lower bounds on the former. This shows that the maintenance of $Q$ in the insert-delete setting can be more expensive than in the insert-only setting.
@@
Third, we devise new dynamic query evaluation algorithms that use our reductions. We call them MVIVM (short for MultiVariate IVM). They improve or recover best known update times, e.g., amortized constant update time for any acyclic query in the insert-only setting and for any hierarchical query in the insert-delete setting. They also achieve optimal $\widetilde{\mathcal{O}}(|\mathcal{D}|^{1/2})$ amortized update time for Loomis-Whitney queries in the insert-delete setting, where $\mathcal{D}$ is the database size at update time.
@@
Fourth, our algorithms support constant-delay enumeration of both the full query output and the change to the output (delta enumeration) after each update. In particular, they map an input stream of updates to an output stream of updates, so they are closed under updates. In practice, the change to the output after an update is often much smaller than the full output, thus using delta enumeration can be more efficient than full enumeration.
Fifth, we prove more general lower bounds on the update time for any join query $Q$. Prior works, e.g., (Berkholz et al., 2017; Kara et al., 2019), typically rely on fine-grained complexity conjectures, such as the OMv conjecture (Henzinger et al., 2015), to prove such lower bounds and only for specific queries. In contrast, our lower bounds are stated in terms of any lower bounds on the static query evaluation problem, regardless of what these lower bounds are. Hence, our lower bounds can be thought of as unconditional in the sense that they do not rely on any conjecture that might be true or false. They also apply to any query.
1.1. Motivating Examples
@@
$Q_{4}(A,B,C,D)$ $=R(A,B,C)\wedge S(B,C,D)\wedge T(A,C,D)\wedge U(A,B,D)$
We would like to perform single-tuple updates to each query, while ensuring constant-delay enumeration of the tuples in the query output after each update. Theses queries are not $q$-hierarchical, hence they cannot admit worst-case constant time per single-tuple update (Berkholz et al., 2017): The OMv conjecture implies that there is no algorithm that takes $\mathcal{O}(|\mathcal{D}|^{1/2-\gamma})$ time for single-tuple updates, for any $\gamma>0$, to any of these queries. In fact, for a database of size $\mathcal{D}$, systems like DBToaster (Koch et al., 2014), DynYannakakis (Idris et al., 2017), CROWN (Wang et al., 2023), and F-IVM (Nikolic and Olteanu, 2018), need at least $\mathcal{O}(|\mathcal{D}|)$ time for a single-tuple update to each of these queries. For the triangle query $Q_{\triangle}$, IVMϵ takes $\mathcal{O}(|\mathcal{D}|^{1/2})$ amortized time for single-tuple updates (Kara et al., 2020a). In the insert-only setting, CROWN maintains $Q_{3p}$ in amortized constant time per insert. However, it cannot handle non-acyclic queries like $Q_{\triangle}$ and $Q_{4}$ (with update time better than recomputation from scratch). Appendix A explains how each prior approach maintains $Q_{3p}$ and $Q_{\triangle}$ in the insert-delete setting, and compares their update times to our approach, MVIVM.
We show that, for $Q_{3p}$, MVIVM performs single-tuple updates: in amortized constant time in the insert-only setting; and in amortized $\widetilde{\mathcal{O}}(|\mathcal{D}|^{1/2})$ time in the insert-delete setting ($\tilde{O}$ notation hides a $\text{\sf polylog}(|\mathcal{D}|)$ factor). Both times are optimal, the latter is conditioned on the OMv conjecture. The amortized constant time per insert matches that of CROWN (Wang et al., 2023) and does not contradict the non-constant lower bound for non-hierarchical queries without self-joins (Berkholz et al., 2017): that lower bound proof relies on a reduction from OMv that requires both inserts and deletes. Achieving this amortized constant time, while also allowing constant-delay enumeration after each insert, is non-trivial. The lazy maintenance approach, which only updates the input data, is not sufficient to achieve this result. This is because, while it does indeed take constant insert time, it still requires the computation of the query output before the enumeration of the first output tuple. This computation takes linear time using a factorized approach (Kara et al., 2023b) and needs to be repeated for every enumeration request.
@@ 2. Problem Setting
We consider natural join queries, or equivalently, full conjunctive queries and refer to them as queries. Sec. 4 defines them.
We investigate the incremental view maintenance (IVM) of a query $Q$ under single-tuple updates to an initially empty database $\mathcal{D}$. A single-tuple update is an insert or a delete of a tuple into a relation in $\mathcal{D}$. We consider four variants of the IVM problem depending on the update and enumeration mode. With regard to updates, we distinguish between the insert-only setting, where we only allow single-tuple inserts, and the insert-delete setting, where we allow both single-tuple inserts and single-tuple deletes. With regard to enumeration, we distinguish between the full enumeration and the delta enumeration, where, after each update, the full query result and respectively the change to the query result at the current time can be enumerated with constant delay. Sec. 4 introduces the notion of constant-delay enumeration. The four IVM variants are parameterized by a query $Q$ and take as input a database $\mathcal{D}$, whose relations are initially empty, and a stream $\delta_{1}\mathcal{D},\delta_{2}\mathcal{D},\ldots,\delta_{N}\mathcal{D}$ of $N$ single-tuple updates to $\mathcal{D}$. Neither $N$ nor the updates are known in advance. The updates arrive one by one. Let $\mathcal{D}^{(0)}\stackrel{{\scriptstyle\text{def}}}{{=}}\emptyset,\mathcal{D}^{(1)},\ldots,\mathcal{D}^{(N)}$ be the sequence of database versions, where for each $\tau\in[N]$, the database version $\mathcal{D}^{(\tau)}$ results from applying the update $\delta_{\tau}\mathcal{D}$ to the previous version $\mathcal{D}^{(\tau-1)}$.
@@
The task of the second IVM variant is to support constant-delay enumeration of the change to the query result after each update:
@@
The change $\delta_{\tau}Q(\mathcal{D})$ is the difference between the query result right after update $\tau$ and the query result right after update $\tau-1$.
@@
Consider the triangle query $Q_{\triangle}$ from Eq. (1). Suppose we have the update stream of length $8$ given in the second column of Table 1. The third and fourth column show the full result and the delta result, respectively, after each update.

### v4

@@ Abstract.
We study the dynamic query evaluation problem: Given a full conjunctive query $Q$ and a sequence of updates to the input database, we construct a data structure that supports constant-delay enumeration of the tuples in the query output after each update.
We show that a sequence of $N$ insert-only updates to an initially empty database can be executed in total time $\mathcal{O}(N^{\mathsf{w}(Q)})$, where $\mathsf{w}(Q)$ is the fractional hypertree width of $Q$. This matches the complexity of the static query evaluation problem for $Q$ and a database of size $N$. One corollary is that the amortized time per single-tuple insert is constant for $\alpha$-acyclic full conjunctive queries.
In contrast, we show that a sequence of $N$ inserts and deletes can be executed in total time $\widetilde{\mathcal{O}}(N^{\mathsf{w}(\widehat{Q})})$, where $\widehat{Q}$ is obtained from $Q$ by extending every relational atom with extra variables that represent the “lifespans” of tuples in the database. We show that this reduction is optimal in the sense that the static evaluation runtime of $\widehat{Q}$ provides a lower bound on the total update time for the output of $Q$. Our approach achieves amortized optimal update times for the hierarchical and Loomis-Whitney join queries.
@@ 1. Introduction
Answering queries under updates to the database, also called dynamic query evaluation, is a fundamental problem in data management, with recent work addressing it from both systems and theoretical perspectives. There have been recent efforts on building systems for dynamic query evaluation, such as DBToaster (Koch et al., 2014), DynYannakakis (Idris et al., 2017, 2020), F-IVM (Nikolic and Olteanu, 2018; Nikolic et al., 2020; Kara et al., 2024), and CROWN (Wang et al., 2023). By allowing tuples to carry payloads with elements from rings (Koch, 2010), such systems can maintain complex analytics over database queries, such as linear algebra computation (Nikolic et al., 2014), collection programming (Koch et al., 2016), and machine learning models (Kara et al., 2024). There has also been work on dynamic computation for intersection joins (Tao and Yi, 2022) and for expressive languages such as Datalog (Motik et al., 2019), Differential Datalog (Abadi et al., 2015), and DBSP (Budiu et al., 2023). The descriptive complexity of various recursive Datalog queries under updates, such as reachability, has also been investigated (Datta et al., 2018; Schwentick et al., 2020). Dynamic query evaluation has also been utilized for complex event recognition (Giatrakos et al., 2017; Grez and Riveros, 2020).
@@
While this problem attracted continuous interest in academia and industry over the past decades, it is only relatively recently that the first results on the fine-grained complexity and in particular on the optimality of this problem have emerged. These efforts aim to mirror the breakthrough made by the introduction of worst-case optimal join algorithms (Atserias et al., 2013; Veldhuizen, 2014; Ngo et al., 2018). Beyond notable yet limited explorations, understanding the optimality of maintenance for the entire language of conjunctive queries remains open. Prime examples of progress towards the optimality of query maintenance are the characterizations of queries that admit (worst-case or amortized) constant time per single-tuple update (insert or delete): the $q$-hierarchical queries (Berkholz et al., 2017; Idris et al., 2017; Kara et al., 2024; Wang et al., 2023), queries that become $q$-hierarchical in the presence of free access patterns (Kara et al., 2023a) or by rewriting under functional dependencies (Kara et al., 2024), or queries on update sequences whose enclosureness is bounded by a constant (Wang and Yi, 2020). The $\delta_{1}$-hierarchical queries (Kara et al., 2020b, 2023b, 2023a) and the triangle queries (Kara et al., 2019, 2020a) admit optimal, albeit non-constant, update time conditioned on the Online Matrix-Vector Multiplication (OMv) conjecture (Henzinger et al., 2015).
In this paper, we introduce a new approach to incremental view maintenance for full conjunctive queries (or queries for short). Our approach complements prior work in four ways.
@@
First, we give reductions from the dynamic query evaluation problem for a query $Q$ to the static query evaluation problem of a derived query $\widehat{Q}$ (called the multivariate extension of $Q$), and vice versa. Our reductions link the complexities of the two problems and allow to transfer both algorithms and lower bounds from one problem to the other. Specifically, we give a reduction from dynamic query evaluation to a wide class of algorithms for static query evaluation, namely algorithms that meet the fractional hypertree width, thus allowing us to use those algorithms for dynamic query evaluation. Moreover, we give a reduction from static query evaluation to any algorithm for dynamic query evaluation, thus allowing us to use lower bounds on static query evaluation to infer lower bounds on dynamic query evaluation. Both reductions use a translation of the time dimension in the dynamic problem into a spatial dimension in the static problem, so that the maintenance under a stream of updates corresponds to taking the intersection of intervals representing the “lifespans” of tuples in the update stream.
@@
Second, we devise new dynamic query evaluation algorithms that use our reductions. We call them MVIVM (short for MultiVariate IVM). Their single-tuple update times are amortized. For any $\alpha$-acyclic query in the insert-only setting, MVIVM take amortized constant single-tuple update time. This recovers a prior result (Wang et al., 2023), when restricted from free-connex to full conjunctive queries. For Loomis-Whitney queries in the insert-delete setting, MVIVM take amortized $\widetilde{\mathcal{O}}(|\mathcal{D}|^{1/2})$ update time, where $|\mathcal{D}|$ is the database size at update time; this is optimal up to a polylogarithmic factor in the database size. This recovers prior work on the triangle query, which is the Loomis-Whitney query with three variables (Kara et al., 2020a). For hierarchical queries in the insert-delete setting, MVIVM take amortized constant single-tuple update time; this is weaker than prior work (Berkholz et al., 2017), which showed that worst-case constant single-tuple update time can be achieved for $q$-hierarchical queries.
@@
Third, MVIVM support (worst-case, not amortized) constant delay enumeration of the tuples in the query output (full enumeration) or in the change to the query output (delta enumeration) after each update. In particular, MVIVM map an input stream of updates to an output stream of updates, so they are closed under updates. In practice, the change to the output after an update is often much smaller than the full output, making delta enumeration more efficient than full enumeration.
Fourth, we pinpoint the complexity gap between the insert-only and the insert-delete settings. In the insert-only setting, we show that one can readily use a worst-case optimal join algorithm like GenericJoin (Ngo et al., 2018) and LeapFrogTrieJoin (Veldhuizen, 2014) to execute a stream of inserts one insert at a time in the same overall time as if the entire stream was executed as one bulk update (which corresponds to static evaluation). This implies an upper bound on the amortized update time for the dynamic evaluation of $Q$ in the insert-only setting. In contrast, in the insert-delete setting, a stream of both inserts and deletes can be executed one update at a time in the same overall time as the static evaluation of a query $\widehat{Q}$, which is obtained by extending every relational atom in $Q$ with extra variables representing the “lifespans” of tuples. This query $\widehat{Q}$ may have a higher complexity than $Q$, thus leading to an upper bound on the amortized update time for $Q$ in the insert-delete setting that may be higher than in the insert-only setting. Nevertheless, our reduction from the maintenance of $Q$ in the insert-delete setting to the static evaluation problem for $\widehat{Q}$ is optimal: Lower bounds on the latter problem imply lower bounds on the former. This shows that maintenance in the insert-delete setting can be more expensive than in the insert-only setting.
Motivating Examples
@@
$Q_{4}(A,B,C,D)$ $=R(B,C,D)\wedge S(A,C,D)\wedge T(A,B,D)\wedge U(A,B,C)$
We would like to perform single-tuple updates to each query, while ensuring constant-delay enumeration of the tuples in the query output after each update. These queries are not $q$-hierarchical, hence they cannot admit worst-case constant time per single-tuple update (Berkholz et al., 2017): The OMv conjecture implies that there is no algorithm that takes $\mathcal{O}(|\mathcal{D}|^{1/2-\gamma})$ time for single-tuple updates, for any $\gamma>0$, to any of these queries. In fact, for a database $\mathcal{D}$, systems like DBToaster (Koch et al., 2014), DynYannakakis (Idris et al., 2017), F-IVM (Nikolic and Olteanu, 2018), and CROWN (Wang et al., 2023) need at least worst-case $\mathcal{O}(|\mathcal{D}|)$ time per single-tuple update to each of these queries. For the triangle query $Q_{\triangle}$, IVMϵ takes $\mathcal{O}(|\mathcal{D}|^{1/2})$ amortized time for single-tuple updates (Kara et al., 2020a). In the insert-only setting, CROWN maintains $Q_{3p}$ in amortized constant time per insert. However, it cannot handle non-acyclic queries like $Q_{\triangle}$ and $Q_{4}$ (with update time better than recomputation from scratch). Appendix A explains how each prior approach maintains $Q_{3p}$ and $Q_{\triangle}$ in the insert-delete setting, and compares their update times to our approach, MVIVM.
We show that, for $Q_{3p}$, MVIVM performs single-tuple updates: in amortized constant time in the insert-only setting; and in amortized $\widetilde{\mathcal{O}}(|\mathcal{D}|^{1/2})$ time in the insert-delete setting ($\tilde{O}$ notation hides a $\text{\sf polylog}(|\mathcal{D}|)$ factor). Both times are optimal, the latter is conditioned on the OMv conjecture. The amortized constant time per insert matches that of CROWN (Wang et al., 2023) and does not contradict the non-constant lower bound for non-hierarchical queries without self-joins (Berkholz et al., 2017): that lower bound proof relies on a reduction from OMv that requires both inserts and deletes. Achieving this amortized constant time, while also allowing constant-delay enumeration after each insert, is non-trivial. The lazy maintenance approach, which only updates the input data, is not sufficient to achieve this result. This is because, while it does indeed take constant insert time, it still requires the computation of the query output before the enumeration of the first output tuple. This computation takes linear time using a factorized approach (Kara et al., 2024) and needs to be repeated for every enumeration request.
@@
2. Preliminaries
In this section, we review some preliminaries, most of which are standard in database theory.
Data Model
Following standard terminology, a relation is a finite set of tuples and a database is a finite set of relations (Abiteboul et al., 1995). The size $|R|$ of a relation $R$ is the number of its tuples. The size $|\mathcal{D}|$ of a database $\mathcal{D}$ is given by the sum of the sizes of its relations.
Queries
(2) $Q(\bm{X})=R_{1}(\bm{X}_{1})\wedge\cdots\wedge R_{k}(\bm{X}_{k}),$
where each $R_{i}$ is a relation symbol, each $\bm{X}_{i}$ is a tuple of variables, and $\bm{X}=\bigcup_{i\in[k]}\bm{X}_{i}$. We refer to $\bm{X}_{i}$ as the schema of $R_{i}$ and treat schemas and sets of variables interchangeably, assuming a fixed ordering of variables. We call each $R_{i}(\bm{X}_{i})$ an atom of $Q$. We denote by $\text{\sf at}(Q)=\{R_{i}(\bm{X}_{i})\mid i\in[k]\}$ the set of all atoms of $Q$ and by $\text{\sf at}(X)$ the set of atoms $R_{i}(\bm{X}_{i})$ with $X\in\bm{X}_{i}$. The set of all variables of $Q$ is denoted by $\text{\sf vars}(Q)=\bigcup_{i\in[k]}\bm{X}_{i}$. Since the set $\bm{X}$ is always equal to $\text{\sf vars}(Q)$ for all queries we consider in the paper, we drop $\bm{X}$ from the query notation and abbreviate $Q(\bm{X})$ with $Q$ in Eq. (2). We say that $Q$ is without self-joins if every relation symbol appears in at most one atom.
The domain of a variable $X$ is denoted by $\text{\sf Dom}(X)$. A value tuple $\bm{t}$ over a schema $\bm{X}=(X_{1},\ldots,X_{n})$ is an element from $\text{\sf Dom}(\bm{X})\stackrel{{\scriptstyle\text{def}}}{{=}}\text{\sf Dom}(X_{1})\times\cdots\times\text{\sf Dom}(X_{n})$. We denote by $\bm{t}(X)$ the $X$-value of $\bm{t}$.
Query Classes
A query is ($\alpha$-)acyclic if we can construct a tree, called join tree, such that the nodes of the tree are the atoms of the query (coverage) and for each variable, it holds: if the variable appears in two atoms, then it appears in all atoms on the path connecting the two atoms (connectivity(Brault-Baron, 2016).
Single-tuple Updates and Deltas
Constant-Delay Enumeration
Width Measures
[…]
Every feasible solution $(\lambda_{R(\bm{X})})_{R(\bm{X})\in\text{\sf at}(Q)}$ to the above program is called a fractional edge cover of $Q$ with respect to $\mathcal{D}$. The optimal objective value of the program is called the fractional edge cover number $\rho^{*}(Q,\mathcal{D})$ of $Q$ with respect to $\mathcal{D}$. The fractional edge cover number gives an upper bound on the size of $Q(\mathcal{D})$: $|Q(\mathcal{D})|\leq 2^{\rho^{*}(Q,\mathcal{D})}$ (Atserias et al., 2013). This bound is known as the AGM-bound and denoted as $\text{\sf AGM}(Q,D)$. We denote by $\rho^{\ast}(Q)$ the optimal objective value of the linear program that results from the above program by replacing the objective function with $\sum_{R(\bm{X})\in\,\text{\sf at}(Q)}\lambda_{R(\bm{X})}$.
Definition 2.1 (Tree Decomposition).
  1. (1)
  2. (2)
Definition 2.2 (The restriction $Q_{\bm{Y}}$ of a query $Q$).
Definition 2.3 (Fractional Hypertree Width).
(5) $\mathsf{w}(\mathcal{T},\chi)\stackrel{{\scriptstyle\text{def}}}{{=}}\max_{t\in V(\mathcal{T})}\rho^{*}(Q_{\chi(t)}),\quad\quad\mathsf{w}(Q)\stackrel{{\scriptstyle\text{def}}}{{=}}\min_{(\mathcal{T},\chi)\in\text{\sf TD}(Q)}\mathsf{w}(\mathcal{T},\chi)$
We call a tree decomposition $(\mathcal{T},\chi)$ of $Q$ optimal if $\mathsf{w}(\mathcal{T},\chi)=\mathsf{w}(Q)$. The fractional hypertree width of a union $Q=\bigcup_{i\in[n]}Q_{i}$ of join queries is $\mathsf{w}(Q)\stackrel{{\scriptstyle\text{def}}}{{=}}\max_{i\in[n]}\mathsf{w}(Q_{i})$.
If a query is a union of join queries, its AGM-bound (or fractional hypertree width) is the maximum AGM-bound (or fractional hypertree width) over its components.
Computational Model
We consider the RAM model of computation where schemas and data values are stored in registers of logarithmic size and operations on them can be done in constant time. We assume that each materialized relation with schema $\bm{X}$ is implemented by a data structure of size $\mathcal{O}(|R|)$ that can: (1) look up, insert, and delete tuples in constant time, and (2) enumerate the tuples in $R$ with constant delay. Given $\bm{Y}\subseteq X$ and $\bm{t}\in\text{\sf Dom}(\bm{Y})$, the data structure can: (1) check $\bm{t}\in\pi_{\bm{Y}}R$ in constant time; and (2) enumerate the tuples in $\sigma_{\bm{Y}=\bm{t}}R$ with constant delay.
@@ 3. Problem Setting
We consider natural join queries, or equivalently, full conjunctive queries and refer to them as queries. We investigate the incremental view maintenance (IVM) of a query $Q$ under single-tuple updates to an initially empty database $\mathcal{D}$. A single-tuple update is an insert or a delete of a tuple into a relation in $\mathcal{D}$. We consider four variants of the IVM problem depending on the update and enumeration mode. With regard to updates, we distinguish between the insert-only setting, where we only allow single-tuple inserts, and the insert-delete setting, where we allow both single-tuple inserts and single-tuple deletes. With regard to enumeration, we distinguish between the full enumeration and the delta enumeration, where, after each update, the full query result and respectively the change to the query result at the current time can be enumerated with constant delay. The four IVM variants are parameterized by a query $Q$ and take as input a database $\mathcal{D}$, whose relations are initially empty, and a stream $\delta_{1}\mathcal{D},\delta_{2}\mathcal{D},\ldots,\delta_{N}\mathcal{D}$ of $N$ single-tuple updates to $\mathcal{D}$. Neither $N$ nor the updates are known in advance. The updates arrive one by one. Let $\mathcal{D}^{(0)}\stackrel{{\scriptstyle\text{def}}}{{=}}\emptyset,\mathcal{D}^{(1)},\ldots,\mathcal{D}^{(N)}$ be the sequence of database versions, where for each $\tau\in[N]$, the database version $\mathcal{D}^{(\tau)}$ results from applying the update $\delta_{\tau}\mathcal{D}$ to the previous version $\mathcal{D}^{(\tau-1)}$.
@@
The task of the second IVM variant is to support constant-delay enumeration of the change to the query result after each update, given by Eq. (3):
@@
Consider the triangle query $Q_{\triangle}$ from Eq. (1). Suppose we have the update stream of length $8$ given in the second column of Table 1. The third and fourth column show the full result and the delta result, respectively, after each update. We will use this example as a running example throughout the paper.